OpenStudy (kpop4life123):

If (7^2)^x = 1, what is the value of x? Explain your answer.

OpenStudy (anonymous):

anything raised to the 0 power is 1

OpenStudy (kpop4life123):

i dont understand

OpenStudy (anonymous):

that's just a property of exponents. 3^0 = 1 49^0 = 1 (ax² - 5)^0 = 1 If you get 1 as the result, (and the base isn't 1), then the exponent has to be 1

OpenStudy (kpop4life123):

so youre saying 0 = 1 ?

OpenStudy (anonymous):

no. I'm saying that x = 0.

OpenStudy (anonymous):



OpenStudy (kpop4life123):

ohhhhhhhhhhhhhhhhh ok